The area of the triangle is 36 square inches. The base is half the height. Find the height and base of the triangle.
Answer by cleomenius(959)

Let x = base
Let 2x = height
36 = 1/2 ( x)(2x)
36 = x^2
x = 6 in
2x = 12 in.
36 = 1/2( 6) (12)
This does check.
Cleomenius.